Attempt Solve for x. 405-5x Express the answer to the hundredths place (i.e., two digits after the decimal point) x=11 140.30 3 9
Answer #1

given,
405 = 5*x^3
x^3 = 405/5
x^3 = 81
x = (81)^1/3
= 4.32675
= 4.33

Answer: 4.33